OpenStudy (anonymous):

b^2/3 b^2/3 ______________ -b^1/3 I understand most of how to do this, but what I don't understand is what to do with the negative be and how to do it.

jhonyy9 (jhonyy9):

in numerator because there is the same terms on exponent you need just add those exponents so hence there will be

jhonyy9 (jhonyy9):

\[b ^{4/3}\]

jhonyy9 (jhonyy9):

in denominator is cube roote from -b

jhonyy9 (jhonyy9):

\[\sqrt[3]{-b}\]

jhonyy9 (jhonyy9):

\[b ^{4/3}/(-b)^{1/3}\]

jhonyy9 (jhonyy9):

when you divide the same terms with exponent you need just subtract the exponents so in this case 4/3 - 1/3 = 3/3 =1 so in this case will get b on exponent 1 duvided by (-1)

jhonyy9 (jhonyy9):

\[b ^{4/3}/(-(b)^{1/3})= b ^{(4/3 - 1/3)}/(-1)=b ^{3/3}/(-1)= b/(-1)=-b\]